### v1

# UniRec: Unified Multimodal Encoding for LLM-Based Recommendations

Abstract
Large language models (LLMs) have recently shown promise for multimodal recommendation, particularly with text and image inputs. Yet real-world recommendation signals extends far beyond these modalities. To reflect this, we formalize recommendation features into four modalities: text, images, categorical features, and numerical attributes, and emphasize unique challenges this heterogeneity poses for LLMs in understanding multimodal information. In particular, these challenges arise not only across modalities but also within them, as attributes (e.g., price, rating, time) may all be numeric yet carry distinct meanings. Beyond this intra-modality ambiguity, another major challenge is the nested structure of recommendation signals, where user histories are sequences of items, each carrying multiple attributes. To address these challenges, we propose UniRec, a unified multimodal encoder for LLM-based recommendation. UniRec first employs modality-specific encoders to produce consistent embeddings across heterogeneous signals. It then applies a triplet representation—comprising attribute name, type, and value—to separate schema from raw inputs and preserve semantic distinctions. Finally, a hierarchical Q-Former models the nested structure of user interactions while maintaining their layered organization. On multiple real-world benchmarks, UniRec outperforms state-of-the-art multimodal and LLM-based recommenders by up to 15%, while extensive ablation studies further validate the contributions of each component.
ulab-uiuc/UniRec[Uncaptioned image] Hugging Face Collection
1 Introduction
Large language models (LLMs) have recently transformed recommender systems by reframing recommendation as a language modeling task (Geng et al., 2022; Li et al., 2023a; Zhang and others, 2023; Bao et al., 2023). Leveraging world knowledge and reasoning ability, LLM-based recommenders can capture rich semantic representations of users and items, enabling zero-shot prediction and explainable recommendations (Hou and others, 2023; Wang et al., 2025). However, most existing approaches primarily operate in text-centric settings, or at best combine text with images, where descriptive content such as reviews, metadata, or product visuals is abundant. While text and visual modalities are important, real-world recommendation data is far more heterogeneous, encompassing numerical, categorical, temporal, and geographical attributes that current LLM-based systems are ill-equipped to handle. Therefore, an open challenge remains: how can we design a unified framework that enables LLMs to effectively understand and reason over heterogeneous multimodal recommendation signals?
Addressing this challenge requires encoders that can faithfully represent such heterogeneous data. An effective encoder must be schema-aware, distinguishing attributes like price versus timestamp even when both are numeric; hierarchy-aware, capturing the nested structure of user histories as sequences of items with multiple attributes; and modality-aware, balancing signals across text, images, categorical fields, and numerical values. Naïve serialization into text or simple concatenation of embeddings discards these structural cues, obscures cross-feature interactions, and fails to capture sequential or relational patterns (Zhou et al., 2023a; Hou and others, 2023; Liu and others, 2023; Singh and others, 2023; Bao et al., 2023).
While recent studies have begun exploring multimodal foundation models for recommendation (Geng et al., 2023; Luo et al., 2024), and vision–language models such as BLIP-2 demonstrate how Q-Formers can bridge modalities (Li et al., 2023b), these approaches are still limited in not tailoring to specific modality pairs or tasks and not providing a general solution for arbitrary heterogeneous recommendation inputs. A unified framework is still missing—one that can preserve schema, modality, and structural information while making heterogeneous signals accessible to LLM reasoning.
To fill this gap, we propose UniRec, a unified multimodal encoder that enables LLMs to leverage heterogeneous recommendation signals. UniRec employs modality-specific encoders to produce aligned embeddings for text, image, categorical, and numerical features. Each attribute is represented as a triplet—(attribute name, type, value)—to disentangle schema from raw inputs and preserve semantic distinctions. A hierarchical Query-Former then aggregates these representations, first into item-level embeddings and subsequently into user-level histories, explicitly maintaining the layered structure of interactions. Integrated with a pretrained LLM, UniRec enables reasoning over multimodal signals without losing structural context, thereby improving recommendation accuracy.
We validate UniRec on a suite of benchmark datasets spanning diverse recommendation scenarios with differing combinations of textual, visual, numerical, categorical, temporal, and geographical attributes. Across all settings, UniRec consistently outperforms state-of-the-art multimodal and LLM-based baselines and demonstrates robustness across datasets with varying attribute compositions, underscoring its effectiveness in both conventional and richly multimodal recommendation tasks.
2 Preliminaries
In this section, we first introduce the preliminaries of multimodal recommendation and formally define the problem setup. We then discuss the limitations of existing methods in this setting, which motivates the design of our proposed UniRec encoder.
2.1 Problem Setup
We begin by formalizing the task of multimodal sequential recommendation, where users interact with items over time and items are described by heterogeneous attributes across multiple modalities.
Users and Items. Let $\mathcal{U}=\{u_{1},\dots,u_{|\mathcal{U}|}\}$ be the set of users and $\mathcal{I}=\{i_{1},\dots,i_{|\mathcal{I}|}\}$ the set of items. Each user $u\in\mathcal{U}$ generates a chronological interaction history
$\mathcal{H}_{u}=\big[(t_{1},\ell_{1},i_{1}),(t_{2},\ell_{2},i_{2}),\dots,(t_{T},\ell_{T},i_{T})\big],$
where $t_{j}$ is the timestamp, $\ell_{j}$ an optional location, and $i_{j}\in\mathcal{I}$ the item at step $j$. The next-item prediction task is to model the conditional distribution
$p(i_{T+1}\mid\mathcal{H}_{u}),$
and use it to rank candidate items for recommendation.
Attributes and Modalities. Each item $i\in\mathcal{I}$ is described by a heterogeneous collection of attributes spanning multiple modalities, such as textual descriptions, images, categorical labels, numerical values, or spatiotemporal information. Formally, let $\mathcal{N}$ denote the attribute namespace and $\{\mathcal{V}_{a}\}_{a\in\mathcal{N}}$ the corresponding value domains. Then an item is associated with
$\mathcal{A}(i)=\{(a,v)\mid a\in\mathcal{N},\,v\in\mathcal{V}_{a}\},$
where $a$ is the attribute name and $v$ its observed value. Different attributes may share the same value format (e.g., both price and rating are numerical) yet carry distinct semantics.
2.2 Limitations of Existing Methods
Although prior work has made progress by leveraging textual and visual signals, existing approaches still face fundamental shortcomings when applied to heterogeneous multimodal recommendation.
Loss of Schema and Structural Semantics. Text-only LLM-based recommenders (Geng et al., 2022; Li et al., 2023a; Ye et al., 2024) flatten multimodal signals into text prompts. This process erases distinctions between different attribute roles and obscures the nested structure of interactions, making it difficult to preserve schema semantics or accurately model structured data.
Shallow Fusion of Modalities. Conventional multimodal recommenders (He and McAuley, 2016a; Wei et al., 2019b; Tao et al., 2022) typically combine auxiliary signals such as images or reviews through simple concatenation or late fusion. Such shallow integration limits the model’s ability to capture fine-grained cross-modal dependencies or hierarchical structures within user–item interactions.
Limited Generalization Across Modalities. Recent LLM–multimodal hybrids (Geng et al., 2023; Luo et al., 2024; Zhang et al., 2025a; López-Ávila and Du, 2025) demonstrate improved multimodal reasoning but are usually tailored to specific modality pairs and lack explicit schema-awareness. This reduces their generalizability across diverse recommendation settings with heterogeneous attribute types.
3 UniRec: Unified Multimodal Encoding for LLM-Based Recommendations
To address the limitations identified in the previous section, we propose UniRec for multimodal LLM-based recommendation. We first formalize heterogeneous signals into four modalities with modality-specific encoders to obtain reliable feature representations. On top of these, we design a schema-preserving triplet representation and a hierarchical Q-Former aggregation mechanism, enabling LLMs to effectively understand heterogeneous recommendation signals for next-item prediction. The details are introduced as follows.
Figure 1: UniRec Model Architecture: (a) Item-Level Q-Former: Raw item attributes across heterogeneous modalities (text, categorical, image, numerical) are processed by modality-specific encoders and triplet formation. These generate schema-aware attribute embeddings, which are then aggregated by the Item Q-Former to produce a fixed-length item representation ($\mathbf{z}_{t}$). (b) User-Level Q-Former: A user’s chronological interaction history, consisting of learned item tokens ($\mathbf{z}_{t}$), multimodal review contexts ($\mathbf{c}_{t}$), and timestamp embeddings ($\mathbf{p}_{t}$), is processed by an Interaction Feature Assembly module. The resulting sequence of combined interaction embeddings is then distilled by the User Q-Former into a unified user representation ($\mathbf{U}$). The arrow passing the Learned Item Tokens from (a) to (b) explicitly models the nested structure of recommendation signals—where a user’s history is a sequence of items, and each item is a collection of heterogeneous attributes. (c) LLM-Based Recommendation: The learned user representation and item representations are projected as soft prompts to condition the LLM for next-item prediction, ranking against a corpus of candidate item embeddings.
3.1 Modality-Specific Encoders
Robust modality-wise encoders and dense, comparable representations are crucial for stable multimodal fusion (Li et al., 2024; Vouitsis et al., 2024; Xu et al., 2022). To this end, we map all inputs into 1024-dimensional embeddings using modality-specific encoders. For text (e.g., titles, reviews), we employ Qwen3-0.6B Embedding [^1] https://huggingface.co/Qwen/Qwen3-Embedding-0.6B. Categorical labels (e.g., product categories) are encoded with the same model using category-aware instructions, avoiding the mixing of sparse one-hot features with dense vectors—a practice known to cause semantic misalignment and training instability (Guan and others, 2022; Li and others, 2022; Cheng and others, 2022). For images, we use CLIP ViT-L/14 (Radford et al., 2021b) [^2] https://huggingface.co/openai/clip-vit-large-patch14, followed by a projection layer that maps the native 768D representations to 1024D. For numerical features, we adopt a Fourier-based Math-Aware Number Encoder (Zhou et al., 2025; Cao et al., 2025), which integrates Fourier components (sine/cosine with log-spaced frequencies), raw magnitude/sign values, and a small learned projection. The encoder is trained with objectives enforcing additivity, invertibility, and distance preservation, while domain-specific adaptations handle temporal cycles (e.g., hour-of-day, month-of-year) and geospatial coordinates projected onto the unit sphere. Implementation details are provided in Appendix A.
3.2 Hierarchical Q-Former Encoder
To model the nested structure of user-item interactions, we introduce a two-stage Hierarchical Q-Former. This architecture first distills an item’s raw multimodal attributes into a fixed-size item representation, and then aggregates a sequence of item interactions into a final user representation.
Schema-Aware Attribute Representation. To preserve the semantic meaning of each attribute (e.g., knowing that ”19.99” is a ”price”), we represent it as a triplet of its $(name,type,value)$. We obtain embeddings for the attribute’s name ($\mathbf{a}_{j}$), its modality type ($\mathbf{t}_{j}$), and its value ($\mathbf{v}_{j}$). These are fused via summation into a single schema-aware attribute embedding $\mathbf{h}_{j}$:
$\mathbf{h}_{j}=\mathbf{a}_{j}+\mathbf{t}_{j}+\mathbf{v}_{j}$
The complete set of attribute embeddings for an item $i$ is denoted by $\mathbf{H}_{i}=\{\mathbf{h}_{1},\dots,\mathbf{h}_{N_{i}}\}$, which serves as the input for the first stage of our hierarchy.
Two-Stage Hierarchical Aggregation. Our aggregation process uses two sequential Q-Formers.
First, an Item Q-Former processes the variable-length set of an item’s attribute embeddings, $\mathbf{H}_{i_{t}}$. Using a set of learnable queries $\mathbf{Q}_{\text{item}}$, it distills this information into a fixed-length item representation, $\mathbf{z}_{t}$:
$\mathbf{z}_{t}=\text{QFormer}_{\text{item}}(\mathbf{Q}_{\text{item}},\mathbf{H}_{i_{t}})\in\mathbb{R}^{K_{\text{item}}\times d}$
Next, a User Q-Former aggregates the sequence of interactions over time. For each step $t$, the item representation $\mathbf{z}_{t}$ is combined with its associated review context embedding $\mathbf{c}_{t}$ and a timestamp embedding $\mathbf{p}_{t}$ to preserve chronological order. A new set of queries, $\mathbf{Q}_{\text{user}}$, processes this entire sequence to yield the final user representation $\mathbf{U}$:
$\mathbf{U}=\text{QFormer}_{\text{user}}(\mathbf{Q}_{\text{user}},\{\text{Concat}(\mathbf{z}_{t},\mathbf{c}_{t})+\mathbf{p}_{t}\}_{t=1}^{T})\in\mathbb{R}^{K_{\text{user}}\times d}$
This design yields rich, multi-token representations for both items ($\mathbf{z}_{t}$) and the user ($\mathbf{U}$), capturing more granular information than a single aggregated vector.
3.3 UniRec Training and Inference
Our training strategy decouples representation learning from LLM adaptation in a two-stage process. First, we pretrain the UniRec encoder with a frozen LLM to learn aligned representations. Second, we fine-tune the encoder and the LLM jointly for the next-item prediction task.
Pretraining Stage. The objective of this stage is to train the modality-specific encoders and the Hierarchical Q-Former to produce a well-structured latent space while the LLM remains frozen. We employ a multi-task learning framework that combines two objectives. The first is a reconstruction loss ($\mathcal{L}_{\text{recon}}$), which ensures the Q-Former’s output retains modality-specific details by using an MLP head to reconstruct the original attribute embeddings from the output tokens. The second is a contrastive loss ($\mathcal{L}_{\text{contrast}}$), which uses InfoNCE (van den Oord et al., 2018) to learn semantic item similarity by treating adjacent items in user histories as positive pairs. The final pretraining objective is a weighted sum of these two losses:
$\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{contrast}}+\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}$
where $\lambda_{\text{recon}}$ is a balancing hyperparameter. This approach ensures the encoder learns representations that are both comprehensive and semantically aligned.
Fine-Tuning Stage. In the second stage, we adapt the system for recommendation by jointly training the Hierarchical Q-Former and the LLM’s Low-Rank Adaptation (LoRA) weights (Hu et al., 2021), while keeping the core modality encoders frozen. The final user representation is projected into the LLM’s word embedding space, functioning as a soft prompt that conditions the LLM on the user’s multimodal history.
The fine-tuning objective is the InfoNCE loss, applied to the next-item prediction task. The model learns to distinguish the ground-truth next item from a set of in-batch negative samples:
$\mathcal{L}_{\text{finetune}}=-\log\frac{\exp(\text{sim}(\mathbf{u},\mathbf{z}_{T+1})/\tau)}{\exp(\text{sim}(\mathbf{u},\mathbf{z}_{T+1})/\tau)+\sum_{j=1}^{N}\exp(\text{sim}(\mathbf{u},\mathbf{z}_{j}^{-})/\tau)}$
This joint training allows the Q-Former to refine its representations for the recommendation task while teaching the LLM to interpret the injected soft prompts.
Inference and Ranking. During inference, a final user representation $\mathbf{u}$ is generated by processing the user’s interaction history and mean-pooling the LLM’s final hidden state. To produce a ranked list, this user embedding is used to compute a relevance score via dot product similarity against a corpus of pre-computed item embeddings, $s(u,i)=\mathbf{u}\cdot\mathbf{z}_{i}$. Candidate items are then ranked in descending order of this score.
4 Experiments
In this section, we conduct a comprehensive set of experiments on diverse and heterogeneous recommendation datasets to evaluate the performance of UniRec, and perform detailed ablation studies to validate the contribution of each design component.
4.1 Experiment Setup
Tasks and Datasets. Table 1 summarizes the attributes available in the datasets used for evaluation. We consider three benchmarks: the Beauty and Baby categories from the Amazon Product Reviews corpus (McAuley et al., 2015), and the Yelp Review dataset (Yelp Inc., 2018). Following common practice, we apply 5-core filtering to retain only users and items with at least five interactions. Each training sample is constructed by extracting 20 consecutive interactions as the historical sequence, with the 21st interaction designated as the ground-truth item. For evaluation, we adopt the leave-one-out strategy: the held-out ground-truth item is ranked against 99 randomly sampled negatives, yielding a candidate set of 100 items per user.
Table 1: Detailed summarization of user and item attributes in Beauty, Baby and Yelp datasets.
Dataset Level Attributes
Beauty User timestamp, rating, title, text, image
Item main_category, title, average_rating, features, description, price, image, store, categories, details
Baby User timestamp, rating, title, text, image
Item main_category, title, average_rating, features, description, price, image, store, categories, details
Yelp User review_date, review_text, review_star
Item name, latitude, longitude, stars, review_count, attributes, categories, image, image_caption, image_label
As shown in Table 1, user-side interactions include timestamps, ratings, and review text, with the Amazon datasets also containing review images. Item-side information covers textual descriptions, categorical labels, numerical values, and images. Amazon data thus provides multimodal signals at both user and item levels, combining product metadata with user-contributed visuals, while Yelp emphasizes spatiotemporal and categorical attributes, such as latitude–longitude coordinates, business categories, and review counts, alongside textual reviews and star ratings. This organization reflects the nested structure of recommendation signals, where each user history is a sequence of items, and each item is described by multiple heterogeneous attributes.
Baselines and Metrics. We evaluate a variety of baseline methods across three scenarios, grouped into (a) Feature-based sequential recommenders, (b) Multimodal recommendation models, and (c) LLM-based multimodal recommenders. For evaluation, we follow the next-item prediction setup with leave-one-out strategy: for each user, the ground-truth item is ranked against 99 randomly sampled negatives. Performance is reported using three standard top-$K$ ranking metrics: Mean Reciprocal Rank (MRR) (Voorhees and Tice, 1999; Cremonesi et al., 2010), Hit Rate at 10 (Hit@10), and Normalized Discounted Cumulative Gain at 10 (NDCG@10) (Järvelin and Kekäläinen, 2002; Burges et al., 2005; Liu, 2009), where higher values indicate better recommendation quality.
  • Multimodal-Feature Sequential Models. This group adapts classical sequential recommenders by enriching item embeddings with multimodal features. Specifically, GRU4Rec (Hidasi et al., 2016), BERT4Rec (Sun et al., 2019), and SASRec (Kang and McAuley, 2018) are implemented using representations extracted from a pre-trained CLIP ViT-B/32 model (Radford et al., 2021a). CLIP encodes product text and images into a shared embedding space, which replaces ID embeddings while leaving the original model architectures unchanged.
  • Multimodal Recommendation Models. These approaches are explicitly designed to integrate multimodal signals into recommendation. VBPR (He and McAuley, 2016a) augments matrix factorization with visual preference vectors derived from product images. MMGCN (Wei et al., 2019b) builds modality-specific user–item graphs and aggregates them with graph neural networks. BM3 (Zhou et al., 2023b) introduces a self-supervised learning framework that applies dropout to generate contrastive views without explicit negative sampling, improving robustness. LGMRec (Guo et al., 2024) decouples collaborative filtering and content modeling by combining local modality-specific graphs with global hypergraph-based embeddings, achieving stronger alignment between content and interaction signals.
  • LLM-based Multimodal Recommenders. Recent methods leverage large language models to reason over multimodal content. IISAN (Fu et al., 2024) adopts a parameter-efficient fine-tuning framework that decouples intra- and inter-modal adaptation, improving both training efficiency and scalability while maintaining accuracy. MLLM-MSR (Ye and others, 2025) takes a summarization approach: it converts product images into textual descriptions, uses an LLM to summarize user histories, and fine-tunes the model for sequential recommendation. These models highlight the emerging trend of directly coupling multimodal inputs with LLM.
Table 2: Performance comparison on 3 datasets using MRR, Hit@10, and NDCG@10. Baseline models are grouped into three categories: multimodal-feature sequential, multimodal recommendation, and LLM-based multimodal. Bold and underline denote best and second-best results.
Beauty Baby Yelp
Model MRR Hit@10 NDCG@10 MRR Hit@10 NDCG@10 MRR Hit@10 NDCG@10
Multimodal-Feature Sequential Models
GRU4Rec 0.2087 0.4215 0.2478 0.1365 0.2780 0.1532 0.4120 0.7815 0.4920
BERT4Rec 0.2215 0.4391 0.2605 0.1422 0.2919 0.1605 0.4182 0.7890 0.4975
SASRec 0.2549 0.4598 0.2890 0.1610 0.3387 0.1860 0.4335 0.8012 0.5128
Multimodal Recommendation Models
VBPR 0.2476 0.4503 0.2818 0.1469 0.3413 0.1773 0.4605 0.8390 0.5433
MMGCN 0.3130 0.5045 0.3465 0.1848 0.3849 0.2170 0.5222 0.8769 0.6018
BM3 0.3305 0.5628 0.3864 0.2166 0.4365 0.2501 0.5405 0.8912 0.6187
LGMRec 0.3433 0.5861 0.4025 0.2279 0.4522 0.2668 0.5530 0.9025 0.6261
LLM-based Multimodal Recommenders
IISAN 0.2513 0.4492 0.2851 0.1476 0.3398 0.1784 0.4627 0.8426 0.5462
MLLM-MSR 0.3249 0.5713 0.3976 0.2084 0.4341 0.2495 0.5561 0.9014 0.6352
UniRec 0.3737 0.6270 0.4449 0.2635 0.4673 0.2977 0.5622 0.9150 0.6489
Implementation Details. We implement UniRec on top of the Qwen3-Embedding-0.6B model with LoRA adaptation and integrated Q-Formers. All modalities are mapped into a unified 1024-dimensional space. Training updates Q-Former and LoRA parameters using the AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, learning rate $1\times 10^{-4}$, weight decay $0.01$) with linear warm-up (20 steps) followed by cosine decay. We employ the InfoNCE loss with temperature 0.07, training for 50 epochs with batch size 16 (accumulation step 1), and evaluate with batch size 32. LoRA is applied to attention and feed-forward layers with rank 16, $\alpha=32$, and dropout 0.1. To stabilize and accelerate training, we enable FP16 precision, gradient clipping (norm 1.0), gradient checkpointing, and sequence length grouping. All experiments are conducted on a single NVIDIA A6000 GPU.
4.2 UniRec Outperforms State-of-the-Art Baselines
In the Beauty, Baby, and Yelp datasets, we compare UniRec against multimodal-feature sequential models, multimodal recommendation models, and recent LLM-based multimodal recommenders, as shown in Table 2. Across all three datasets, UniRec consistently outperforms the strongest baselines by relative margins of up to 16% in MRR, establishing a new state-of-the-art.
In the product recommendation setting on Beauty and Baby, where specialized multimodal models such as LGMRec dominate, UniRec delivers notable improvements—8.8% on Beauty and 15.6% on Baby in MRR—highlighting that its schema- and hierarchy-aware design provides clear advantages in modeling complex multimodal attributes. In the point-of-interest recommendation setting on Yelp, which emphasizes spatiotemporal and categorical diversity, UniRec further surpasses the strongest LLM-based competitor MLLM-MSR with a 1.1% gain in MRR, demonstrating robustness beyond product domains. Together, these results confirm that UniRec generalizes effectively across heterogeneous recommendation scenarios, ranging from multimodal product domains to geographic and categorical POI recommendation, consistently outperforming both classical multimodal architectures and modern LLM-based methods.
Figure 2: Both schema- and hierarchy-aware components are crucial for UniRec ’s performance. Results are shown on Beauty, Baby, and Yelp datasets (measured in MRR). Performance improves step by step as components are added: starting from the minimal configuration (w/o Both), introducing either triplet representation or user-level tokens yields clear gains, while combining both achieves the highest performance.
4.3 Ablation Studies Validate the Design of UniRec
We perform a comprehensive set of ablation studies to validate the effectiveness of UniRec ’s design. By systematically removing or varying its core components, we isolate their individual contributions and assess how each choice impacts the overall performance.
Schema- and Hierarchy-Aware Design Matters. This ablation examines the importance of UniRec ’s schema- and hierarchy-aware components: the triplet representation at the item level and the user-level summarization tokens at the sequence level. By selectively removing these elements, we isolate their individual and joint contributions to recommendation performance.
  • w/o Both. Neither triplet representation nor user-level tokens are used, representing the minimal configuration.
  • w/o Triplet. Removes triplet decomposition of item attributes, while retaining user-level tokens.
  • w/o User Token. Keeps triplet representation for items but discards user-level summarization tokens.
As shown in Figure 2, performance improves consistently as each component is introduced. Starting from the minimal setting (w/o Both), adding either triplet representation or user-level tokens yields notable gains, showing that each contributes independently. The best performance is achieved when both are enabled, confirming their complementary roles: triplet representation structures multimodal item attributes effectively, while user-level tokens capture nested dependencies across interaction histories. Together, these design choices allow UniRec to better exploit schema- and hierarchy-rich recommendation data.
Query-Based Fusion Outperforms Alternatives. We ablate the fusion mechanism used to combine modality-specific embeddings, with results reported in Table 3. All variants adopt the same Qwen3-0.6B embedding LLM as backbone and share identical candidate sets, modality-specific encoders, and training schedules. The only difference lies in the fusion design, for which we compare the following:
  • Pure Text. Uses only item textual descriptions (e.g., titles/reviews) as content features, discarding other modalities. This text-only setup is a common baseline in multimodal recommendation studies (Zhou et al., 2023a; Liu et al., 2023).
  • MLP Fusion. Concatenates embeddings from all modalities into a single vector and processes them with a multilayer perceptron (“early fusion”), a simple but shallow cross-modal strategy (Baltrušaitis et al., 2019).
  • CLIP-Style Projection. Maps each modality into a shared latent space through modality-specific linear layers, with alignment guided by a contrastive objective, following dual-encoder vision–language paradigms such as CLIP and ALIGN (Radford et al., 2021b; Jia et al., 2021).
  • Self-Attention Fusion. Treats modality embeddings as tokens and applies a Transformer-style self-attention layer to capture pairwise interactions (Vaswani et al., 2017).
The results show that Pure Text already forms a strong baseline, highlighting the informativeness of textual signals. MLP Fusion slightly underperforms, suggesting that naive concatenation introduces noise without modeling schema distinctions. CLIP-Style Projection offers modest gains by better aligning text and image, yet its design is limited to pairwise modality alignment and struggles with heterogeneous attributes. Self-Attention Fusion achieves stronger results, confirming the value of richer cross-modal interactions, but still fails to capture hierarchical user–item structures. In contrast, UniRec consistently surpasses all alternatives across datasets and metrics, demonstrating that schema-aware, hierarchical query-based fusion provides a principled and robust solution for multimodal recommendation.
Table 3: Query-based hierarchical fusion achieves the strongest results. Ablation on fusion mechanisms across Beauty, Baby, and Yelp datasets. All models share identical encoders, training schedules, and candidate sets; only the fusion strategy differs. Bold and underline denote the best and second-best results.
Beauty Baby Yelp
Fusion Mechanism MRR Hit@10 NDCG@10 MRR Hit@10 NDCG@10 MRR Hit@10 NDCG@10
Pure Text 0.3025 0.5016 0.3358 0.1857 0.3941 0.2179 0.4870 0.8643 0.5710
MLP 0.2980 0.4950 0.3301 0.1830 0.3865 0.2135 0.4825 0.8570 0.5655
CLIP 0.3150 0.5205 0.3482 0.1915 0.4040 0.2230 0.4962 0.8725 0.5795
Self-attention 0.3330 0.5635 0.3920 0.2140 0.4365 0.2520 0.5288 0.8955 0.6120
UniRec 0.3737 0.6270 0.4449 0.2635 0.4673 0.2977 0.5622 0.9150 0.6489
Token Count Sensitivity Reveals a Sweet Spot. We examine how the number of latent tokens in the item-level and user-level Q-Formers affects performance (Figure 3). The results show that token count is crucial for balancing expressiveness and generalization. At the item level, accuracy peaks at 4 tokens, after which additional tokens yield diminishing or even negative returns, suggesting that a compact set of latent tokens suffices to capture key multimodal attributes. In contrast, the user-level Q-Former benefits from a slightly larger capacity, with optimal performance around 4–8 tokens, reflecting the greater complexity of modeling long interaction histories.
Figure 3: Optimal token counts emerge for both item- and user-level Q-Formers. Left: item-level tokens. Right: user-level tokens. Each curve shows MRR on one dataset, and the red star ($\star$) marks the token count achieving the highest performance.
Overall, the trends reveal a trade-off: too few tokens underfit the data, while too many introduce redundancy, overfitting, or unstable training. These findings highlight the importance of careful token calibration at both item and user levels to ensure robust hierarchical representation learning.
5 Related Work
Multimodal Recommendation. A large body of research has shown that incorporating auxiliary modalities such as text, images, or reviews can substantially enrich user and item representations and alleviate data sparsity (He and McAuley, 2016b; Wei et al., 2019a; Tao and others, 2022; Zhou et al., 2023a; Liu and others, 2023; Yu et al., 2025). These works established that multimodal signals carry complementary semantics beyond ID-based interactions and can improve personalization in various domains. More recent efforts have begun to align multimodal content with pretrained language models, as in VIP5 (Geng et al., 2023), demonstrating the promise of combining vision, language, and recommendation. Collectively, these advances highlight multimodality as a powerful driver for next-generation recommender systems.
LLM-based Recommendation. The rise of large language models has introduced a new paradigm where recommendation is formulated as a language modeling problem (Geng et al., 2022; Li et al., 2023a; Zhang and others, 2023; Bao et al., 2023; Hou and others, 2023). By unifying diverse tasks into a text-to-text format, LLM-based recommenders benefit from pretrained world knowledge, zero-shot generalization, and explainability. Recent surveys further consolidate their versatility across domains (Hou and others, 2023; Wang et al., 2025). In parallel, multimodal LLMs such as BLIP-2 (Li et al., 2023b) illustrate how vision–language pretraining enables cross-modal reasoning, suggesting new opportunities for recommendation scenarios that span heterogeneous signals. These developments collectively point to unifying multimodal content with LLM reasoning as a natural and promising next step.
6 Conclusion
We presented UniRec, a unified multimodal encoder that models heterogeneous user–item–attribute signals through schema-aware triplet representations and a hierarchical Q-Former, enabling LLMs to reason effectively for recommendation. Experiments on multiple benchmarks showed consistent state-of-the-art performance with notable improvements over multimodal and LLM-based baselines, while ablations confirmed the value of hierarchical fusion and compact tokenization. Although UniRec depends on clean attribute schemas and was tested primarily in offline next-item prediction, it opens promising directions for incorporating richer modalities, handling schema noise, adapting to online and continual settings, and addressing fairness and efficiency concerns. We hope this framework inspires future progress toward general-purpose and trustworthy multimodal recommendation systems.
Ethics Statement
All authors of this paper have read and adhered to the ICLR Code of Ethics. Our work does not involve human subjects, personal data, or sensitive attributes. We followed best practices for data usage, ensured compliance with licensing terms, and considered potential risks of bias or misuse.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our results. Details of the model architecture, training settings, and hyperparameters are described in Section 4. All datasets we used are publicly available. The training scripts and evaluation code will be released upon publication to facilitate replication.
Appendix A Encoder Implementation Details
Text Encoder
We employ the Qwen3-0.6B embedding model (Zhang et al., 2025b), an instruction-tuned encoder. For categorical inputs, we prepend descriptors (e.g., “Category:”), which stabilizes training and prevents semantic drift between categorical fields and free-form textual descriptions.
Image Encoder
Images are processed with CLIP ViT-L/14 (Radford et al., 2021b), which produces 512-dimensional embeddings aligned with text through large-scale contrastive pretraining. This backbone offers a strong balance between representational quality and computational efficiency, enabling reliable multimodal alignment.
Numerical Encoder
We design a Mathematical-Aware Numerical Encoder inspired by recent work on Fourier- and wavelet-based numerical embeddings (Zhou et al., 2025; Cao et al., 2025). Given a scalar input, the Mathematical-Aware Numerical Encoder generates a high-dimensional embedding using:
  • Fourier features: sine and cosine components with logarithmically spaced frequencies, capturing periodic structure across scales;
  • Raw value features: two dimensions encoding magnitude and sign, preserving linear ordering;
  • Learned projection: residual nonlinear features for task-specific representational capacity.
The encoder is trained with a multi-objective loss that enforces:
  1. 1.
    Additivity: $E(a+b)\approx E(a)+E(b)$, encouraging arithmetic structure in the embedding space;
  2. 2.
    Invertibility: a small decoder reconstructs the original scalar, ensuring information preservation;
  3. 3.
    Distance preservation: triplet loss enforces that embedding distances reflect numeric differences.
Normalization is performed conservatively with bounded scaling factors to maintain these mathematical properties during training.
Domain-Specific Numerical Features
Temporal features are decomposed into secular and cyclical components. Secular time normalizes absolute timestamps, while cyclical features (hour-of-day, day-of-week, day-of-year, month-of-year) are encoded with sine/cosine functions, ensuring continuity across cycle boundaries (e.g., 23:59 and 00:01 map to nearby embeddings).
Geospatial coordinates are projected to the unit sphere and represented in 3D Cartesian coordinates ($x=\cos(\text{lat})\cos(\text{lon}),y=\cos(\text{lat})\sin(\text{lon}),z=\sin(\text{lat})$). This preserves great-circle distances, which are more faithful to real-world geography than raw latitude/longitude. A small neural projection refines these representations for downstream use.
Appendix B Detailed Rationale for Triplet Representation and Q-Former Interaction
B.1 Limitations of Naive Serialization
A primary limitation of many existing LLM-based recommenders is the loss of schema semantics, which occurs when heterogeneous features are naively serialized into a single text string. For example, an item might be represented as ”Title: Running Shoes, Brand: Nike, Price: 99.99, Rating: 4.5”. While readable to humans, this format obscures the crucial distinction between attributes that may share a data type but carry vastly different meanings. The numerical value ”99.99” for a price has a different semantic role and scale than the value ”4.5” for a rating. When tokenized by an LLM, these distinctions are often lost, forcing the model to re-learn fundamental schema concepts from unstructured text, which is inefficient and error-prone. Our triplet representation avoids this by explicitly disentangling the name, type, and value of each attribute before they are embedded, preserving this critical structural information.
B.2 Mechanism of Schema-Aware Attention in Q-Former
The triplet-based design is foundational for the subsequent hierarchical aggregation performed by the Q-Former. The core of the Q-Former is a cross-attention mechanism, which can be formulated as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$
In our Item Q-Former, the learnable queries $\mathbf{Q}$ are fixed, while the keys $\mathbf{K}$ and values $\mathbf{V}$ are linear projections of the input attribute representations $\{\mathbf{h}_{j}\}_{j=1}^{N_{i}}$.
If the input were an unstructured collection of embeddings from naive concatenation, the keys and values would lack the necessary semantic cues for the queries to perform meaningful summarization. The model would struggle to differentiate an embedding for ”price” from an embedding for ”rating”.
However, by using our triplet-based representation $\mathbf{h}_{j}=\mathbf{e}_{a_{j}}+\mathbf{e}_{t_{j}}+\mathbf{e}_{v_{j}}$, we provide a structured and disentangled input space. Each vector $\mathbf{h}_{j}$ explicitly encodes the attribute’s name, type, and value. Consequently, the keys $\mathbf{K}$ derived from these vectors are rich with schema information. This allows the learnable queries in $\mathbf{Q}$ to specialize. For instance, one query might learn to assign high attention scores to keys corresponding to ”price” attributes, effectively becoming a ”price expert” that extracts cost-related information from items. Another query might specialize in ”brand” or ”category” information. This specialization is only possible because the triplet representation provides the semantic scaffolding necessary for the Q-Former to effectively identify, prioritize, and aggregate information based on its role and content, rather than just its raw value.
Appendix C Conceptual Generalization and Configuration of the Q-Former
C.1 From Modality Bridge to Hierarchical Summarizer
The Querying Transformer (Q-Former), as introduced in vision-language models like BLIP-2 (Li et al., 2023b), was originally conceived as a lightweight bridge between two powerful, frozen encoders (e.g., vision and language). It functions as an information bottleneck, using a small, fixed set of learnable query vectors to extract a fixed-size representation from one modality (vision) that is most relevant to the other (language).
Our work presents a significant conceptual generalization of this role. Instead of bridging two different modalities, we repurpose the Q-Former as a versatile primitive for hierarchical data summarization within the single, complex domain of recommendation. This is achieved through a nested application:
  • Item Q-Former as a Heterogeneous Set Aggregator: At the lower level, the Item Q-Former operates on an unordered set of heterogeneous attribute representations for a single item. Its function is many-to-one aggregation, learning to distill the most salient features from a variable collection of multimodal attributes into a single, canonical item embedding.
  • User Q-Former as a Sequential Aggregator: At the higher level, the User Q-Former performs a more traditional sequence modeling task. It takes the time-ordered sequence of item embeddings and summarizes them into a single user representation, capturing temporal dynamics and evolving preferences.
This hierarchical application showcases that the query-based attention mechanism is a flexible and effective tool for learning to summarize complex data structures, extending its utility far beyond its initial vision-language application.
C.2 Tuning Q-Former Capacity via Token Count
A key design choice in our hierarchical architecture is the number of learnable queries, $K_{\text{item}}$ and $K_{\text{user}}$, used in the Item and User Q-Formers, respectively. These values directly determine the number of output latent tokens and thus control the capacity of the information bottleneck at each level of the hierarchy. This token count is a critical hyperparameter that balances representational expressiveness and model complexity.
A small number of tokens forces the model to learn a highly compressed, dense representation, which may be efficient but could fail to capture the full richness of the input data (underfitting). Conversely, a large number of tokens increases the model’s capacity but may introduce redundancy, capture noise, and increase the risk of overfitting, in addition to raising computational costs. The optimal token count may also differ between the item and user levels, given that one summarizes a set of static attributes while the other summarizes a dynamic sequence of interactions. As such, we conduct a detailed ablation study in our experiments (Section  3) to identify the optimal ”sweet spot” for both $K_{\text{item}}$ and $K_{\text{user}}$, ensuring robust and effective representation learning.
Appendix D Benefits of Decoupled Pretraining
Our two-stage training strategy is critical for the model’s success. Attempting to train the entire system end-to-end from scratch would require the LLM to simultaneously learn to interpret noisy, unaligned multimodal signals while also mastering the recommendation task. This process is both computationally prohibitive and prone to instability.
By first pretraining the UniRec encoder with a carefully designed multi-objective loss, we create a well-structured latent space where user and item embeddings are meaningfully aligned before the LLM is engaged. The reconstruction loss ensures that the Q-Former’s compressed representations do not discard vital information from any modality, while the contrastive loss organizes the latent space according to semantic similarity. This provides the LLM with clean, ”pre-digested,” and semantically rich inputs during the fine-tuning stage, simplifying its adaptation task. This paradigm of decoupling representation learning from generative fine-tuning mirrors the effective training strategies of state-of-the-art vision-language models like BLIP-2 (Li et al., 2023b).
Appendix E Dataset Details
Below is the detailed statistics for Dataset used in our training and evaluation, the number of users, items and sparsity is reported in 4, while a sample of each dataset is reported in 5, 6,  7 and  8
Table 4: Statistics of the benchmark datasets used for the next-item prediction task.
Dataset # Users # Items Sparsity
Beauty 22,363 12,101 99.9267%
Baby 19,445 7,050 99.8827%
Yelp-2018 77,278 45,639 99.9403%
Table 5: An example of a multimodal item from the Amazon dataset, showcasing the variety of attributes available for a single product.
Attribute Value
parent_asin B07G9GWFSM
title Lurrose 100Pcs Full Cover Fake Toenails Artificial Transparent Nail Tips Nail Art for DIY
main_category All Beauty
store Lurrose
average_rating 3.7
rating_number 35
price \$6.99
features The false toenails are durable with perfect length. You have the option to wear them long or clip them short, easy to trim and file them to in any length and shape you like. ABS is kind of green enviromental material, and makes the nails durable, breathable, light even no pressure on your own nails. Fit well to your natural toenails. Non toxic, no smell, no harm to your health. Wonderful as gift for girlfriend, family and friends. The easiest and most efficient way to do your toenail tips for manicures or nail art designs…
description Description: The false toenails are durable with perfect length… Plus, ABS is kind of green enviromental material… Feature: - Color: As Shown.- Material: ABS.- Size: 14.3 x 7.2 x 1cm. Package Including: 100 x Pieces fake toenails
details Color: As Shown Size: Large Material: Acrylonitrile Butadiene Styrene (ABS) Brand: Lurrose Style: French Product Dimensions: 5.63 x 2.83 x 0.39 inches; 1.9 Ounces UPC: 799768026253 Manufacturer: Lurrose
images Present (2 images, MAIN variant shown)
Table 6: An example of a user interaction from the Amazon dataset. This includes the user’s review, rating, and associated metadata for a specific item.
Attribute Value
user_id AEYORY2AVPMCPDV57CE337YU5LXA
parent_asin B08BBQ29N5
asin B088SZDGXG
sort_timestamp 1634275259292 (2021-10-15)
rating 3.0 / 5.0
verified_purchase Yes
helpful_votes 0
title Meh
text These were lightweight and soft but much too small for my liking. I would have preferred two of these together to make one loc. For that reason I will not be repurchasing.
images Present (1 image)
Table 7: An example of a multimodal item from the Yelp dataset, showcasing the variety of attributes available for a single business.
Attribute Value
business_id tnhfDv5Il8EaGSXZGiuQGg
name Garaje
address 475 3rd St
city San Francisco
state CA
postal_code 94107
coordinates latitude = 37.7817529521,  longitude = -122.39612197
stars 4.5 / 5.0
review_count 1198
is_open Yes (1)
attributes RestaurantsTakeOut = true BusinessParking: garage = false,  street = true,  validated = false,  lot = false,  valet = false
categories Mexican,  Burgers,  Gastropubs
hours Monday: 10:00–21:00 Tuesday: 10:00–21:00 Wednesday: 10:00–21:00 Thursday: 10:00–21:00 Friday: 10:00–21:00 Saturday: 10:00–21:00 Sunday: 11:00–18:00
photo_id _nN_DhLXkfwEkwPNxne9hw
photo.business_id tnhfDv5Il8EaGSXZGiuQGg
caption carne asada fries
label food
Table 8: An example of a user profile and associated review from the Yelp dataset, illustrating user activity and interaction with a business.
Attribute Value
user_id Ha3iJu77CxlrFm-vQRs_8g
name Sebastien
review_count 56
yelping_since 2011-01-01
friends wqoXYLWmpkEH0YvTmHBsJQ KUXLLiJGrjtSsapmxmpvTA 6e9rJKQC3n0RSKyHLViL-Q
useful (given) 21
review_id zdSx_SD6obEhz9VrW9uAWA
funny (given) 88
cool (given) 15
fans 1032
elite 2012 2013
business_id tnhfDv5Il8EaGSXZGiuQGg
stars 4 / 5
date 2016-03-09
text Great place to hang out after work: the prices are decent, and the ambience is fun. It’s a bit loud, but very lively. The staff is friendly, and the food is good. They have a good selection of drinks.
useful (received) 0
funny (received) 0
cool (received) 0
Appendix F LLM Writing Usage Disclosure
An LLM was utilized solely as a writing assistant to refine grammar and improve sentence readability. Its role was limited to enhancing linguistic clarity, with no involvement in shaping the research design, conducting data analysis, or influencing the interpretation of results.